OpenStudy (anonymous):

what is the missing term in the geometric sequence -3,__,__,-375

OpenStudy (tkhunny):

Two Missing Terms?

OpenStudy (jdoe0001):

hey you, hey thou, us speaking to thee

OpenStudy (anonymous):

hey you, hey thou, us speaking to thee

OpenStudy (tkhunny):

Geometric Sequence. Each term differs from the previous term by a common ratio. 1 1*2 = 2 2*2 = 4 4*2 = 8 1, 2, 4, 8 is a geometric sequence beginning with 1 and having common ratio 2.